# Detect, Explain, Escalate: Low-Carbon Dialogue Breakdown Management for LLM-Powered Agents

### v2

@@ Abstract
While Large Language Models (LLMs) are transforming numerous applications, their susceptibility to conversational breakdowns remains a critical challenge undermining user trust. This paper introduces a “Detect, Explain, Escalate” framework to manage dialogue breakdowns in LLM-powered agents, emphasizing low-carbon operation. Our approach integrates two key strategies: (1) We fine-tune a compact 8B-parameter model, augmented with teacher-generated reasoning traces, which serves as an efficient real-time breakdown ‘detector’ and ‘explainer.’ This model demonstrates robust classification and calibration on English and Japanese dialogues, and generalizes well to the BETOLD dataset, improving accuracy by 7% over its baseline. (2) We systematically evaluate frontier LLMs using advanced prompting (few-shot, chain-of-thought, analogical reasoning) for high-fidelity breakdown assessment. These are integrated into an ‘escalation’ architecture where our efficient detector defers to larger models only when necessary, substantially reducing operational costs and energy consumption. Our fine-tuned model and prompting strategies establish new state-of-the-art results on dialogue breakdown detection benchmarks, outperforming specialized classifiers and significantly narrowing the performance gap to larger proprietary models. The proposed monitor-escalate pipeline reduces inference costs by 54%, offering a scalable, efficient, and more interpretable solution for robust conversational AI in high-impact domains. Code and models will be publicly released.
Index Terms:
large language models, conversational artificial intelligence, human-computer interaction, dialogue systems, model calibration
@@ I Introduction
Conversational artificial intelligence (AI) has experienced exponential growth driven by the rapid advancement and widespread adoption of large language models (LLMs) such as ChatGPT [1], Claude [2], and Llama [3]. The unprecedented success of ChatGPT, which attracted one million users in five days and exceeded 100 million in two months [4, 5], highlights both the unique capabilities and the immense potential of LLMs. These generative models have rapidly permeated numerous sectors, including customer service, content creation, marketing, and education, fundamentally reshaping human-AI interactions [6, 7, 8].
@@
Despite these advancements, the rapid integration of LLMs into critical domains has surfaced significant challenges, notably dialogue breakdowns, which can severely disrupt user trust and conversational effectiveness [9, 10, 11]. Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [12, 13]. Table I presents an example where the assistant’s utterances are contradictory, illustrating a common breakdown scenario.
Addressing dialogue breakdowns becomes increasingly critical with the increasing adoption of LLMs in high-stakes environments [6, 7, 8]. Moreover, the tendency of LLMs to produce overly confident yet potentially inaccurate or hallucinated responses further complicates their reliable deployment [14, 15]. Consequently, there is a pressing need for robust methodologies to detect and mitigate dialogue breakdowns, thereby preserving conversational quality and user trust.
Previous research in dialogue breakdown detection has predominantly focused on specialized classifiers trained on labeled datasets like the Dialogue Breakdown Detection Challenge [12, 16]. However, the generalization of these models to diverse, real-world contexts remains limited. Recent literature also indicates that generalist LLMs, despite their flexibility, still lag behind human-level performance in nuanced conversational tasks, indicating a persistent research gap [17, 18].
Addressing these challenges, we introduce approaches that leverage the reasoning capabilities of generalist LLMs through supervised fine-tuning and structured prompting. Specifically, we fine-tune the efficient and powerful Llama-3.1 8B model [3] on both English and Japanese tracks from the DBDC5 dataset, incorporating synthetic reasoning trajectories generated by a larger, more capable Llama-3.3 70B model. These distilled reasoning tracks, incorporated during fine-tuning, aim to improve the student model’s decision-making process and enhance the interpretability of its predictions. We rigorously evaluate its generalization capability using the challenging BETOLD dataset [19], which explicitly focuses on task-oriented dialogue breakdowns across diverse conversational contexts.
@@
Additionally, we provide a comprehensive comparative analysis of both closed-source frontier models (OpenAI [1] and Anthropic [2]) and open-source alternatives (Meta [3], Mistral AI [20], and DeepSeek [21]), specifically targeting their dialogue breakdown detection capabilities. We explore advanced prompting methodologies including few-shot learning [22, 23], zero-shot chain-of-thought prompting [24], and propose analogical reasoning enhanced by curricular learning strategies [25, 26, 27]. These strategies encourage systematic reasoning and adaptability, often reducing the need for extensive labeled datasets.
Through these novel techniques, we establish state-of-the-art benchmarks for the Dialogue Breakdown Detection Challenge. Furthermore, we examine calibration metrics alongside traditional classification metrics, gaining insights into model overconfidence and reliability. Our proposed architecture introduces a cost-effective, real-time dialogue breakdown monitoring system that utilizes a fine-tuned Llama-3.1 8B model, selectively invoking resource-intensive models such as GPT-4, DeepSeek-R1, Sonnet-3.5, or Llama-3.1 405B when necessary. Our architecture optimizes resource utilization by strategically employing high-capacity models, which typically incur significant operational costs and energy consumption, only when necessary. This approach achieves substantial cost reductions and promotes sustainability.
Our contributions are summarized as follows:
@@
  • Conducted the first comprehensive comparative analysis across a wide range of open-source and frontier closed-source models, establishing new benchmarks in dialogue breakdown detection.
  • Extensively evaluated model calibration alongside accuracy, revealing critical insights into reliability and overconfidence across diverse conversational scenarios.
@@
  • Proposed an efficient, real-time deployment architecture that significantly reduces operational costs and promotes sustainability by selectively invoking large-scale models.
@@ II Related Work
Ensuring robustness in conversational AI systems, particularly in dialogue breakdown detection and mitigation, is a vital area of research. Dialogue breakdowns refer to situations where conversational coherence or relevance is disrupted, hindering the conversation’s smooth progression and diminishing user satisfaction [12, 13]. Effective conversational agents must not only detect these breakdowns but also address them to maintain user engagement and trust [28, 29].
@@ II-A Specialized Models for Dialogue Breakdown Detection
Research into dialogue breakdown detection has yielded specialized classifiers that push the performance on datasets like DBDC5 (see Section V-A for more information) [12]. The top-performing approaches leverage pre-trained transformer [30] encoders fine-tuned for breakdown classification. For example, the best model on the DBDC5 English track was an augmented BERT-based classifier, BERT+SSMBA [31, 32]. BERT+SSMBA incorporates unlabeled dialogue data via two complementary techniques: extended pre-training on dialogue-rich datasets such as Reddit and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [33]. This approach exploits unlabeled data to improve the classifier’s robustness in detecting dialogue breakdowns.
Further advances in semi-supervised learning have also significantly impacted the domain. The S2T2 model, for example, introduced a dual-teacher training paradigm to leverage both labeled and unlabeled dialogue data [34]. S2T2 employs two teacher models: one trained on high-quality labeled data and another trained on masked dialogues variations, to guide a student model collaboratively. This hybrid training strategy has achieved new state-of-the-art results on the DBDC5 dataset, surpassing earlier methods. Specifically, S2T2 utilizes RoBERTa-large for the English track and XLM-R-large (a multilingual variant of RoBERTa) combined with a context-matching mechanism for the Japanese track [35, 36, 34].
Recent studies have also explored the potential of general-purpose large language models (LLMs) in dialogue breakdown detection. Finch et al. [17, 18] evaluated ChatGPT’s performance in identifying dialogue behaviour across nine distinct categories defined in the ABC-Eval dataset. Their findings reveal that while ChatGPT shows promise, even outperforming specialized models in identifying empathetic behaviour, it still falls short of human performance in other dialogue tasks. This suggests that current large models still face challenges in reliably identifying breakdown instances despite their sophisticated conversational abilities and extensive world knowledge. Nonetheless, using LLMs for dialogue evaluation is a promising direction since they bring broad world knowledge and understanding of conversational dynamics. To our knowledge, no other work has evaluated LLMs for dialogue breakdown detection and remediation.
@@ II-B Current State of Conversational Agents
Beyond detecting breakdowns, state-of-the-art conversational agents incorporate strategies to mitigate breakdowns. For example, if a breakdown is detected, a system might employ a recovery strategy (as studied in DBDC5’s recovery track) to ask for clarification or provide corrections [12]. Leading chatbots like ChatGPT and Claude are also trained via alignment techniques such as reinforcement learning from human feedback (RLHF) to minimize toxic, incoherent, or nonsensical outputs that could derail a dialogue [1, 2]. Models like Claude-2 even engage in internal debates during training to identify and eliminate reasoning flaws, aiming to reduce the likelihood of breakdowns due to factual errors or contradictions [37]. The net result is that modern conversational AI systems are progressively improving at maintaining coherent dialogues despite ambiguous user inputs.
@@ II-C Techniques in Conversational AI
To further improve the reliability of LLM conversational agents, researchers are exploring a variety of techniques:
Analogical Reasoning. This technique guides a language model to draw on relevant past examples or scenarios by analogy when confronting a new problem. Instead of relying solely on provided examples, the model generates analogous examples as context. Inspired by human problem-solving via analogy, analogical prompting has been shown to improve reasoning accuracy [25, 38]. It “prompts language models to self-generate relevant exemplars or knowledge in context, before proceeding to solve the given problem,” thereby removing the need for hand-crafted exemplars and tailoring the reasoning to the specific query [25]. In practice, a dialogue agent using analogical reasoning might recall a similar conversation or scenario it has encountered and use that information to form a better response to a user query.
Chain-of-Thought Reasoning. Chain-of-thought (CoT) is a reasoning technique in which the model is encouraged to produce explicit intermediate reasoning steps (“thinking aloud”) before giving a final answer. This has effectively improved logical consistency and arithmetic or commonsense reasoning in LLMs [39, 40, 24]. For dialogue agents, chain-of-thought can be used internally (the model can reason through the user’s query, context, and knowledge base step-by-step) to formulate a correct and context-appropriate response. This reduces mistakes and non-sequiturs. CoT reasoning is now a common technique to bolster LLM robustness on tasks requiring multi-step inference or clarification, as it significantly “enhances their ability to reason” about the conversation [24, 40, 39].
Zero-Shot and Few-Shot Learning. Unlike traditional dialogue systems requiring extensive task-specific training, LLM-based agents excel at in-context learning as they can adapt to new instructions or domains given few examples or none at all. In few-shot learning, a handful of example dialogues or question-answer pairs are provided in the prompt, and the model generalizes the pattern to the new query. The seminal GPT-3 work showed that large models can perform new language tasks by observing just a few demonstrations, essentially treating “language models as few-shot learners” [22]. This allows a conversational agent to be quickly customized to a new domain by giving it a few sample conversations as guidance rather than retraining the model. LLMs can often do this for dialogue tasks. For instance, a well-designed prompt can instruct the model to play the role of a customer service agent and will generate reasonable dialogue without any fine-tuning. In-context learning empowers LLM-based agents to handle various conversational scenarios with minimal additional training.
@@ III Problem Definition
Dialogue breakdown, defined as the deterioration of coherence, relevance, or conversational fluidity between a user and a conversational agent, remains a significant challenge in conversational AI[12, 29]. Such breakdowns may manifest as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that hinder the natural progression of dialogue and consequently degrade user trust and satisfaction [10, 11]. As LLMs from the OpenAI [1], Claude [2], and Llama [3] families are rapidly adopted for diverse conversational tasks, addressing these dialogue disruptions has become paramount. Moreover, dialogue breakdown detection becomes even more critical due to LLMs’ propensity for confident but potentially incorrect or hallucinated outputs, resulting in increased user frustration [14, 15].
We consider a multi-turn dialogue sequence $D$ between a user (U) and a conversational agent (A). The dialogue proceeds in pairs of utterances, where at turn $i$, the user produces an utterance $u_{i}$, and the agent replies with $s_{i}$. Formally, we may represent the dialogue as follows:
@@ III-A Utterance-Level Breakdown Detection
For a complete dialogue $D$ of length $n$ system turns, the output $\mathcal{O}$ of the detection system is thus:
@@ III-B Consolidation of Three-Class Annotations
Previous datasets like DBDC [12] use three-level labels: Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB). Given the subjectivity of human annotations in dialogue breakdown detection, we consolidate ‘Possible Breakdown’ PB into the ‘Breakdown’ (B) class. When multiple human annotators supply these labels, let $p_{i}$ be the fraction who labeled $s_{i}$ as B or PB. We can convert each utterance into a binary label $b_{i}$ by thresholding:
$b_{i}\;=\;\begin{cases}1,&\text{if }p({b_{i}}\mid s_{i})\,\geq\,0.5,\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
where $p({b_{i}}\mid s_{i})$ is the fraction of human annotators labeling $s_{i}$ as a breakdown. Throughout this paper, the binary classification of “breakdown versus non-breakdown” refers to this mapping.
III-C Conversation-Level Labeling
In task-oriented datasets such as BETOLD [19], the focus is on the entire conversation outcome. Each dialogue $D$ is labeled a failure if, for instance, the user hangs up or requests escalation to a human agent. We denote this conversation-level label as

### v3

# Detect, Explain, Escalate: Sustainable Dialogue Breakdown Management for LLM Agents

@@ Abstract
Large Language Models (LLMs) have demonstrated substantial capabilities in conversational AI applications, yet their susceptibility to dialogue breakdowns poses significant challenges to deployment reliability and user trust. This paper introduces a “Detect, Explain, Escalate” framework to manage dialogue breakdowns in LLM-powered agents, emphasizing resource-efficient operation. Our approach integrates two key strategies: (1) We fine-tune a compact 8B-parameter model, augmented with teacher-generated reasoning traces, which serves as an efficient real-time breakdown detector and explainer. This model demonstrates robust classification and calibration on English and Japanese dialogues, and generalizes to the BETOLD dataset, improving accuracy by 7% over its baseline. (2) We systematically evaluate frontier LLMs using advanced prompting (few-shot, chain-of-thought, analogical reasoning) for high-fidelity breakdown assessment. These are integrated into an “escalation” architecture where our efficient detector defers to larger models only when necessary, substantially reducing operational costs and computational overhead. Our fine-tuned model and prompting strategies achieve state-of-the-art performance on DBDC5 and strong results on BETOLD, outperforming specialized classifiers on DBDC5 and narrowing the performance gap to larger proprietary models. The proposed monitor–escalate pipeline reduces inference costs by 54%, providing a cost-effective and interpretable solution for robust conversational AI in high-impact domains. Code and models will be publicly released.
@@ I Introduction
Conversational artificial intelligence (AI) has experienced substantial growth driven by advances in large language models (LLMs) such as ChatGPT [openai_gpt-4_2023], Claude [ClaudeModelCard], and Llama [llama]. ChatGPT’s rapid adoption, reaching 100 million users within two months [Yatoo2023ChatGPT, Khosravi2023Chatbots], demonstrates the widespread interest in LLM-based conversational systems and their potential to transform human-computer interaction. These models are now embedded in applications spanning customer service, content creation, marketing, and education, enabling new forms of human–AI interaction [Currie2023A, Gabriella2024The, Arviani2023ChatGPT].
@@
Despite these advancements, the accelerated integration of LLMs into critical domains has revealed significant challenges in dialogue breakdown detection and mitigation, which directly impact user trust and conversational effectiveness [Ghassel2024Are, Fan2022Building, candello_recovering_2018]. Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [higashinaka_dialogue_2016, Ashktorab]. Table I illustrates a common breakdown pattern, where the assistant’s utterances contradict one another.
As LLM-based systems are deployed in high-stakes environments [Currie2023A, Gabriella2024The, Arviani2023ChatGPT], robust handling of such breakdowns becomes increasingly important. The tendency of LLMs to produce overly confident yet inaccurate or hallucinated outputs further complicates safe deployment [xiong_can_2023, zhou_larger_2024]. There is therefore a pressing need for methods that not only detect dialogue breakdowns reliably but also support mitigation strategies that preserve conversational quality and user trust.
Traditionally, dialogue breakdown detection has been approached using specialized classifiers trained on labelled datasets such as the Dialogue Breakdown Detection Challenge (DBDC) [higashinaka_dialogue_2016, tsunomori-etal-2018-evaluating]. While these models achieve strong benchmark performance, their generalization to diverse, real-world contexts remains limited. At the same time, recent work shows that general-purpose LLMs still lag behind human performance for nuanced conversational behavior, despite their broad coverage and world knowledge [finch_leveraging_2023, finch_dont_2023]. This gap indicates that neither specialized classifiers nor generalist LLMs alone provide a complete solution.
In this paper, we propose a framework that combines the strengths of both approaches. We leverage the reasoning capabilities of generalist LLMs through supervised fine-tuning and structured prompting, and we embed these capabilities within a resource-aware deployment architecture.
First, we fine-tune a parameter-efficient Llama-3.1 8B model [llama] on both English and Japanese tracks of the DBDC5 dataset [Higashinaka2020DBDC5], augmenting supervision with synthetic reasoning trajectories generated by a larger teacher model, Llama-3.3 70B. These distilled reasoning traces are used during training to shape the student model’s decision process and to support interpretable justifications at inference time. We then evaluate the model’s generalization on the BETOLD dataset [terragni2022_BETOLD], which focuses on task-oriented dialogue breakdowns in realistic customer-service settings.
@@
Second, we conduct a comprehensive comparative analysis of both closed-source frontier models (OpenAI [openai_gpt-4_2023], Anthropic [ClaudeModelCard]) and open-source alternatives (Meta [llama], Mistral [jiang_mixtral_2024], DeepSeek [deepseek]) on dialogue breakdown detection. We explore and propose novel prompting schemes, including few-shot learning [brown_language_2020, palm2023], chain-of-thought prompting [wei_chain--thought_2023], and analogical reasoning with curriculum learning [analogical, xu-etal-2020-curriculum, Ryu2024CurricuLLM:], to elicit more systematic reasoning while regulating token costs.
Third, we integrate these components into a hierarchical, cost-aware deployment architecture. A fine-tuned Llama-3.1 8B monitor provides fast, per-turn breakdown detection and explanation and selectively escalates to more capable models (GPT-4, DeepSeek-R1, Claude-3.5 Sonnet, or Llama-3.1 405B) only when its confidence indicates elevated risk. This architecture yields substantial cost and energy savings without sacrificing accuracy.
We summarize our contributions as follows:
@@
  • We conduct, to our knowledge, the first comprehensive comparative analysis of open-source and frontier closed-source LLMs, establishing new benchmarks on dialogue breakdown detection tasks.
  • We jointly evaluate accuracy and calibration, revealing important differences in reliability and overconfidence across models, prompting strategies, and languages.
@@
  • We propose a real-time deployment architecture that significantly reduces operational costs and optimizes resource usage by selectively invoking large-scale models.
  • We demonstrate that monitor-generated explanations can effectively guide a superior model to repair problematic responses; conditioning the superior model on these justifications resolves 97% of sampled breakdowns.
@@ II Related Work
Ensuring robustness in conversational AI systems, particularly for dialogue breakdown detection and mitigation, is a central research challenge. Dialogue breakdowns occur when coherence or relevance is disrupted, hindering the smooth progression of a conversation and reducing user satisfaction [higashinaka_dialogue_2016, Ashktorab]. Practical conversational agents must therefore both detect breakdowns and employ recovery strategies to maintain engagement and trust [mctear, skantze-2017-towards].
@@ II-A Specialized Models for Dialogue Breakdown Detection
Research into dialogue breakdown detection has produced specialized classifiers that perform strongly on benchmarks such as DBDC5 (see Section V-A[higashinaka_dialogue_2016]. Leading approaches rely on fine-tuned transformer encoders [attentionIsAll]. For example, the best-performing model on the DBDC5 English track, BERT+SSMBA [Devlin2019BERT, ng_improving_2020], augments a BERT-based classifier with unlabelled dialogue data through extended pre-training on dialogue-rich corpora (e.g., Reddit) and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [ng_ssmba_2020]. This design improves robustness by exploiting large quantities of unlabelled conversational data.
Semi-supervised approaches further advance the state of the art. The S2T2 model employs a dual-teacher paradigm that leverages both labelled and unlabelled dialogues [lin_semi-supervised_2022]. One teacher is trained on high-quality labelled data, while the other is trained on masked dialogue variants. A student model is jointly guided by both teachers. S2T2 reaches new state-of-the-art performance on DBDC5 using RoBERTa-large on the English track and XLM-R-large with a context-matching mechanism on the Japanese track [liu2019roberta, conneau2020unsupervised, lin_semi-supervised_2022].
More recently, general-purpose LLMs have been evaluated for dialogue breakdown detection. Finch et al. [finch_leveraging_2023, finch_dont_2023] studied ChatGPT’s behaviour across nine categories in the ABC-Eval dataset. While ChatGPT outperforms specialized models on some categories (e.g., empathetic behaviour), it still falls short of human performance on many dialogue tasks. This suggests that current large models, despite their sophisticated capabilities, remain unreliable for fine-grained breakdown detection. To our knowledge, prior work has not systematically evaluated multiple LLM families on dialogue breakdown detection and remediation.
@@ II-B Current State of Conversational Agents
Beyond detection, modern conversational agents increasingly incorporate mechanisms to mitigate or recover from breakdowns. In the DBDC5 recovery track, for instance, systems respond to detected breakdowns by asking clarifying questions or providing corrections [higashinaka_dialogue_2016]. Industrial systems such as ChatGPT and Claude are trained with alignment techniques like reinforcement learning from human feedback (RLHF) to reduce toxic, incoherent, or nonsensical outputs [openai_gpt-4_2023, ClaudeModelCard]. Claude-2, for example, uses internal debate and self-critique during training to identify and minimize reasoning flaws [claude2]. These mechanisms reduce the incidence of breakdowns but do not eliminate them, especially under distribution shift or adversarial inputs.
@@ II-C Techniques in Conversational AI
A variety of techniques have been proposed to improve the reliability of LLM-based conversational agents:
Analogical Reasoning. Analogical prompting encourages a model to draw on relevant past examples by analogy when solving a new problem. Instead of relying on hand-crafted exemplars, the model is prompted to generate analogous examples in context and then solve the target query [analogical, webb_emergent_2023]. This approach can improve reasoning by creating problem-specific exemplars without additional labelling.
Chain-of-Thought Reasoning. Chain-of-thought (CoT) prompting instructs the model to produce intermediate reasoning steps before giving a final answer [zs-cot, fs-cot, wei_chain--thought_2023]. CoT has been shown to improve performance on arithmetic, logical, and commonsense reasoning tasks. For dialogue agents, CoT can be applied internally to reason about user intent, dialogue history, and knowledge bases, thereby reducing non sequiturs and hallucinations.
Zero-Shot and Few-Shot Learning. Unlike traditional dialogue systems requiring extensive task-specific training, LLMs excel at in-context learning, adapting to new tasks given instructions and a small number of examples. In few-shot learning, a prompt includes a handful of example dialogues or question–answer pairs, and the model generalizes the pattern to new inputs. GPT-3 demonstrated that large models can function as few-shot learners without task-specific fine-tuning [brown_language_2020]. For dialogue tasks, this allows rapid adaptation to new domains using carefully crafted prompts rather than retraining.
II-D Hierarchical Architectures
Hierarchical or multi-tiered architectures are a common strategy for balancing performance and cost in AI systems [chen2024frugalgpt, yue2024large]. In the context of LLMs, such architectures typically use smaller, faster models for routine inference and reserve larger, more capable models for complex or high-stakes cases [10.5555/3692070.3693614, 10.5555/3586589.3586709]. This can be viewed as a system-level “mixture of experts” [2017olln], where routing decisions balance accuracy and computational expense. Our work extends this paradigm for real-time dialogue management, positioning a lightweight monitor in front of more capable models to create a practical and sustainable framework for conversational agents.
@@ III Problem Definition
Dialogue breakdown refers to the deterioration of coherence, relevance, or conversational fluency between a user and a conversational agent [higashinaka_dialogue_2016, skantze-2017-towards]. Breakdowns may appear as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that disrupt the dialogue and erode user trust [Fan2022Building, candello_recovering_2018]. As LLMs from the OpenAI [openai_gpt-4_2023], Claude [ClaudeModelCard], and Llama [llama] families are deployed in diverse conversational tasks, detecting and mitigating such breakdowns becomes critical. This is exacerbated by LLMs’ propensity to produce confident but incorrect or hallucinated content [xiong_can_2023, zhou_larger_2024].
We consider a multi-turn dialogue sequence $D$ between a user (U) and a conversational agent (A). At each turn $i$, the user produces an utterance $u_{i}$, and the agent responds with $s_{i}$. The dialogue can be written as:
@@ III-A Utterance-Level Breakdown Detection
For a dialogue $D$ with $n$ system turns, the detector produces:
@@ III-B Consolidation of Three-Class Annotations
Following prior works [Ghassel2024Are, lin_semi-supervised_2022], we preprocess DBDC5’s three-way labels (Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB)) into a binary Breakdown/Non-Breakdown scheme. Ambiguous annotations (PB) are merged into the Breakdown class. When multiple annotators label each system turn, let $p_{i}$ be the fraction who assign B or PB to utterance $s_{i}$. We define a binary label $b_{i}$ by thresholding:
$b_{i}\;=\;\begin{cases}1,&\text{if }p({b_{i}}\mid s_{i})\,\geq\,0.5,\\[4.0pt] 0,&\text{otherwise},\end{cases}$
where $p({b_{i}}\mid s_{i})$ denotes the fraction of annotators labelling $s_{i}$ as a breakdown.
In deployment, the monitor must provide a binary control signal (continue vs. intervene) under tight latency constraints and asymmetric risk: missing a breakdown (false negative) is often more costly than a spurious intervention (false positive). Consolidating PB into the Breakdown class yields a conservative detector that treats boundary cases as risky. This intentionally increases the false-positive rate to reduce false negatives. The decision threshold $T$ on $\hat{c}_{i}$ can then be tuned to balance sensitivity and specificity for a given application.
III-C Conversation-Level Labelling
In task-oriented datasets such as BETOLD [terragni2022_BETOLD], labels are assigned at the conversation level. Each dialogue $D$ is marked as a failure if, for example, the user hangs up or requests escalation to a human agent. We denote this label by: